OpenStudy (anonymous):

Is 3d+4=2+3d-1/2 no solution? Why?

OpenStudy (anonymous):

when you subtract 3d from both sides your variable is going to disappear

OpenStudy (anonymous):

If you group all the d terms, you find that they all drop out, and you are left with 4 = 2 - (1/2) which is FALSE! So no solution

OpenStudy (anonymous):

\[3d+4=2+3d-\frac12\\4=2-\frac12\\4\neq\frac32\]